OpenStudy (nyasia):

A right triangle has side lengths 7, 24 , and 25 as shown below. Use these lengths to find cos a, tan a, and sin a .

OpenStudy (anonymous):

 Which side is which? All I would know is that 25 is the hypotenuse

OpenStudy (anonymous):

Of course, this is also under the impression the triangle is a right

OpenStudy (nyasia):

This is the triangle

OpenStudy (anonymous):

Use the definition of sin, cos, and tanget Sin(A) = Opposite/Hypotenuse Cos(A) = Adjacent/Hypotenuse Tan(A) = Opposite/Adjacent

OpenStudy (anonymous):



OpenStudy (kameron018):

sin= 7/25 cos 24/25 tan= 7/24

OpenStudy (nyasia):

Thank You !!